A tractor of mass 2000kg pulls a trailer of mass 1500kg. The total frictional force is 3000N and acceleration of tractor and trailer is 3ms^-2 calculate:

a. The force exerted on the tractor by the tow bar when acceleration is 3ms^-2

b. The force exerted when did tractor and the trailer are moving on a constant speed of 4ms^-1


1
Expert's answer
2021-08-09T06:51:02-0400

Given:

"M=2000\;\rm kg,\quad m=1500\: kg,"

"\\
 F_f=3000\: {\rm N}."

(a) "a=3\:\rm m/s^2,"

"(M+m)a=F-F_f""F=(M+m)a+F_f\\=(2000+1500)*3+3000=13500\:\rm N"

(b) "a=0,"

"F=F_f=3000\:\rm N"